OpenStudy (anonymous):

@Abhisar A man standing on a street wants to throw a book up to a boy leaning out of a window 5 m above street level.What must be the velocity the man has to throw the book so that it reaches the boy? (g=10ms-2)

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

10m/s , right?

OpenStudy (abhisar):

Sorry i wrote the equation wrong ! \(\color{blue}{\text{Originally Posted by}}\) @Abhisar use the equation \(\large\sf v^2=u^2-2gs\). Here, v=0, s=5. Find the value of u \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

and yes the answer will be 10m/s\(\color{blue}{\text{Originally Posted by}}\) @MichelleShum 10m/s , right? \(\huge\checkmark\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

yes...